OpenStudy (anonymous):

help medal will be given!!!!!! for a 225-watt bulb, the intensity i of light of humens at a distance of x feet is i=225/x^2 A.what is the intensity of light 5ft from the bulb B. suppose your distance from the bulb doubles, how does the intensity of the light change ? explain

OpenStudy (anonymous):

a) l(5) = 225/5^2 = 9 b) Intensity decreases by a factor of 2^2 = 4. So intensity will be 1/4 what it was before distance was doubled Let d = original distance from bulb. Then we get intensity l = 225/d^2 If we double distance (to 2d) we get intensity l = 225/(2d)^2 = 225/(4d^2) = 1/4 (225/d^2) = 1/4 * original intensity

OpenStudy (anonymous):

Hey doll, this guys is right, everything checks out!